OpenStudy (anonymous):

What is the product of 3(sqrt of u) (sqrt448 + 4sqrt7)? Simplify if possible

OpenStudy (anonymous):

\[3\sqrt{u}(\sqrt{448}+4\sqrt{7})\]

OpenStudy (anonymous):